OpenStudy (anonymous):

What is the quotient: (2x^2 + 10x + 12) ÷ (x + 3) ? 2x + 6 2x – 4, r = 1 2x – 6 2x + 4 What is the quotient: (2x^2 + 10x + 12) ÷ (x + 3) ? 2x + 6 2x – 4, r = 1 2x – 6 2x + 4 @Mathematics

OpenStudy (anonymous):

2x+4

OpenStudy (anonymous):

\[\frac{2 x^2+10 x+12}{x+3}=\frac{2 (x+2) (x+3)}{x+3}=2 (x+2)=2x+4 \]

OpenStudy (anonymous):

Thank you